VIBRATIONAL SPECTRA OF CHLOROFORM, FREON-11 AND SELECTED ISOTOPOMERS IN THE TERAHERTZ REGION

\underline{CHRISTA HAASE}, JINJUN LIU^{a} and FRÉDÉRIC MERKT, Laboratorium für Physikalische Chemie, ETH-Zürich, 8093 Zürich, Switzerland.

The fundamental bands of the CCl_{3} asymmetric deformation modes of selected isotopomers of chloroform (CHCl_{3}) and freon-11 (CFCl_{3}) have been measured between 7 and 8 THz in a static cell at ambient temperature using a laser-based source of tunable radiation in the terahertz region (0.1-10 THz) of the electromagnetic spectrum. Simulation of the rotational contour of the ν_{6} fundamental transition of ^{12}CH^{35}Cl_{3} confirmed previously suggested values for C_{6} and C_{6}ζ_{6}.^{b} The fundamental frequencies were derived with a precision of 2 GHz for all compounds except CF^{35}Cl_{3}, where the precision amounted to 3 GHz. The frequencies are in agreement with values calculated ab initio. Extension of the experimental setup to enable measurement of THz spectra of molecules in supersonic jet expansions and partially resolve their rotational structure is in progress.

acurrent address: Department of Chemistry, Laser Spectroscopy Facility, The Ohio State University, 120 W. 18th Avenue, Columbus, Ohio 43210, USA.

bJ. H. Carpenter, P. J. Seo and D. H. Whiffen J. Mol. Spectrosc. \underline{170}, 215, (1995).